OpenStudy (anonymous):

if f(x) = cos2x, find f(3pi/4)

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

f(3/4pi) = cos(3/2Pi) =0

OpenStudy (anonymous):

yes, but how did you figure it out?

OpenStudy (anonymous):

if f(x)=cos(2x) then f(3/4pi) = cos (2*3/4pi) = cos ((3/2pi) But 3/2Pi = 270 degrees ( where Pi = 180) cos (270) = 0

OpenStudy (anonymous):

ohhh okay. thanks. :)

OpenStudy (anonymous):

vw